### v1

@@ Abstract
Molecular structure elucidation from spectra is a foundational problem in chemistry, with profound implications for compound identification, synthesis, and drug development. Traditional methods rely heavily on expert interpretation and lack scalability. Pioneering machine learning methods have introduced retrieval-based strategies, but their reliance on finite libraries limits generalization to novel molecules. Generative models offer a promising alternative, yet most adopt autoregressive SMILES-based architectures that overlook 3D geometry and struggle to integrate diverse spectral modalities. In this work, we present DiffSpectra, a generative framework that directly infers both 2D and 3D molecular structures from multi-modal spectral data using diffusion models. DiffSpectra formulates structure elucidation as a conditional generation process. Its denoising network is parameterized by Diffusion Molecule Transformer, an SE(3)-equivariant architecture that integrates topological and geometric information. Conditioning is provided by SpecFormer, a transformer-based spectral encoder that captures intra- and inter-spectral dependencies from multi-modal spectra. Extensive experiments demonstrate that DiffSpectra achieves high accuracy in structure elucidation, recovering exact structures with 16.01% top-1 accuracy and 96.86% top-20 accuracy through sampling. The model benefits significantly from 3D geometric modeling, SpecFormer pre-training, and multi-modal conditioning. These results highlight the effectiveness of spectrum-conditioned diffusion modeling in addressing the challenge of molecular structure elucidation. To our knowledge, DiffSpectra is the first framework to unify multi-modal spectral reasoning and joint 2D/3D generative modeling for de novo molecular structure elucidation.
@@
Molecular Structure Elucidation, Molecular Spectra, Diffusion Models
@@ 1 Introduction
Molecular structure elucidation represents one of the fundamental challenges in chemistry, materials science, and biology [1, 2, 3], requiring the accurate determination of atomic connectivity, stereochemistry, and three-dimensional (3D) arrangements based on experimental data. Traditional approaches often require expert interpretation and substantial manual effort to propose and validate structural hypotheses. Although these methods have proven invaluable over decades of chemical research, the increasing complexity of natural products, synthetic compounds, and pharmaceutical intermediates demands more sophisticated computational approaches that can efficiently navigate the vast and complex chemical space of possible structures.
@@
The integration of machine learning into spectroscopic analysis has opened new avenues for automating and accelerating molecular structure determination. Early applications of machine learning to molecular structure elucidation primarily adopted retrieval or search-based paradigms [1, 4, 5]. These approaches treat the task as a matching problem: given an experimental spectrum, the model searches within a predefined molecular library to retrieve candidate structures whose spectra best align with the observed one. Although efficient in practice, these methods are inherently constrained by the coverage and quality of the reference library. To mitigate this limitation, some efforts have proposed spectral prediction models, which simulate spectra from known molecular structures [6, 7, 8, 9, 10, 11, 12]. These predicted spectra are then used to expand the searchable library and improve retrieval coverage. However, such methods still rely on the existence of a finite reference library and fundamentally struggle to generalize to novel molecules that are not well represented in the library, limiting their applicability in open-ended or exploratory chemical analyses.
@@
To overcome the reliance on molecular libraries and enable fully data-driven structure elucidation, recent research has turned to more expressive neural architectures capable of de novo structure inference directly from spectra. Initial attempts in this direction framed the task as a predictive problem, training models to infer coarse structural features from spectra, such as molecular properties [13], fingerprints [14], fragments [15, 16], chemical formulas [17], or other approximate structure profiles [18, 19, 20, 21]. Although informative, these coarse predictions do not achieve complete molecular structure elucidation.
@@
In a shift toward complete structure elucidation, subsequent works introduced generative models that take spectral data as input and produce discrete molecular representations, such as SMILES [22] or molecular graphs [23]. These models typically follow autoregressive architectures, generating molecular structures token by token using RNNs [24, 25, 26] or Transformers [27, 16, 4, 28, 29, 30]. Despite their promise, they face two fundamental limitations. First, generative models built on discrete SMILES sequences or graphs inherently lack geometric inductive biases and cannot effectively capture 3D geometric information. This is critical because many spectral modalities, such as infrared (IR), Raman, and ultraviolet-visible (UV-Vis), are inherently dependent on the molecule’s geometry and associated energy states [31]. Second, all existing methods are limited to processing a single type of spectrum in isolation, whereas molecular structure elucidation in practice relies on the joint interpretation of multiple spectral modalities, each offering complementary structural cues [32]. Current models lack support for such multi-modal spectral reasoning, thereby constraining their accuracy and generalizability. These limitations highlight the need for a more principled and unified generative framework that can integrate both geometric reasoning and multi-modal spectral conditioning.
@@
The inherent limitations of existing approaches motivate the development of more sophisticated generative models that can effectively bridge the gap between spectral observations and molecular structure determination. Diffusion models have recently attracted significant attention in generative modeling due to their ability to produce high-quality samples through iterative denoising processes [33, 34]. This framework has demonstrated remarkable success in modeling complex, multi-modal distributions and generating diverse outputs, and has also shown particular promise in molecular applications [35, 3]. The ability of diffusion models to generate chemically meaningful structures and handle multi-modal distributions makes them particularly well-suited for molecular structure elucidation.
@@
In this work, we introduce DiffSpectra, a novel framework that leverages diffusion models to directly generate molecular structures from multi-modal spectra. Our approach addresses the structure elucidation problem as a conditional generation task, where spectral measurements serve as conditioning information to guide the iterative denoising process towards generating plausible molecular structures.
@@
At the core of our framework lies DiffSpectra, a joint 2D and 3D diffusion model that simultaneously generates both molecular graphs and atomic coordinates within a unified diffusion framework. Each molecule is represented as a graph comprising node features (e.g., atom types), edge features (e.g., bond types), and 3D coordinates. This unified representation captures both the topological structure and spatial geometry. DiffSpectra formulates structure elucidation as a continuous-time conditional diffusion process, where noise is progressively added and then removed under the condition of spectral constraints. By simultaneously modeling both topological and geometric components, the model ensures that generated structures satisfy chemical bonding rules while preserving spatial consistency. This dual-space generation is essential for accurately reconstructing molecules whose spectra reflect both bonding connectivities and 3D conformations, effectively leveraging the complementary strengths of 2D topology and 3D geometry to ensure structurally consistent outputs.
@@
To parameterize the denoising process, we introduce the Diffusion Molecule Transformer (DMT), an SE(3)-equivariant architecture designed for molecular structure generation. DMT consists of a stack of transformer-style blocks operating on three interleaved streams—node features, edge features, and coordinates—allowing for rich interaction between topological and geometric representations. Through geometry-aware multi-head attention and equivariant updates, DMT effectively models spatial relationships and maintains chemical symmetry.
@@
To effectively condition the generation process on multi-modal spectra, we employ SpecFormer, a transformer-based spectral encoder specifically designed to process and integrate multiple types of molecular spectra simultaneously, including IR, Raman, and UV-Vis. SpecFormer partitions each spectrum into local patches, embeds them into a unified latent space, and employs multi-head self-attention to capture intra- and inter-spectrum dependencies. To enhance the quality of spectral embeddings, we pre-train SpecFormer using masked patch reconstruction and contrastive alignment with molecular structures. The resulting spectral embeddings provides powerful conditioning signals for DiffSpectra, effectively guiding generation toward candidates that are consistent with the measured spectra.
@@
We extensively evaluate DiffSpectra to assess its effectiveness in structure elucidation from multi-modal spectra. Our results show that DiffSpectra generates chemically valid and stable molecular structures. When conditioned on spectra, our model attains a top-1 exact structure accuracy of 16.01%, and accurately reconstructs key functional groups with over 96% similarity. We further show that using a pre-trained SpecFormer significantly improves structure elucidation accuracy, and that combining multiple spectral modalities yields better performance than using any single spectrum alone. Importantly, we find that sampling multiple candidates greatly enhances the hit rate: the top-20 accuracy rises to 96.86%, indicating that correct structures are consistently ranked among the plausible outputs. This highlights the practical value of diffusion-based sampling in structure elucidation, where generating a shortlist of candidates enables reliable downstream validation. Our ablations further confirm the advantage of model-based SE(3) equivariance and demonstrate that moderate sampling temperatures strike a good balance between diversity and accuracy. These findings highlight the potential of spectrum-guided diffusion models for accurate molecular structure elucidation.
@@ 2.1 Overview of DiffSpectra Framework
To achieve accurate molecular structure elucidation from molecular spectra, we propose DiffSpectra, a spectrum-conditioned diffusion framework that generates molecular structures guided by spectral information. The overall paradigm is illustrated in Figure 1.
@@
Formally, a molecule characterized by its structure and spectra is represented as $\mathcal{M}=(\mathcal{G},\mathcal{S})$. The molecular structure is modeled as a graph $\mathcal{G}=(\mathbf{H},\mathbf{A},\mathbf{X})$, where $\mathbf{H}\in\mathbb{R}^{N\times d_{1}}$ denotes node-level attributes such as atom types and charges, $\mathbf{A}\in\mathbb{R}^{N\times N\times d_{2}}$ encodes pairwise edge features including bond existence and bond types, and $\mathbf{X}\in\mathbb{R}^{N\times 3}$ corresponds to the 3D coordinates of the atoms. Here, $N$ is the number of atoms, $d_{1}$ is the node feature dimensionality, and $d_{2}$ is the edge feature dimensionality. The atomic ordering is consistent across $\mathbf{H}$ and $\mathbf{X}$ to preserve the correspondence between atom types and spatial positions.
@@
Meanwhile, the molecular spectra are denoted as $\mathcal{S}=({\bm{s}}_{1},\ldots,{\bm{s}}_{|\mathcal{S}|})$, where $|\mathcal{S}|=3$ in our setting, covering UV-Vis, IR, and Raman spectra. Concretely, ${\bm{s}}_{1}\in\mathbb{R}^{601}$ records the UV-Vis spectrum from 1.5 to 13.5 eV with 0.02 eV intervals; ${\bm{s}}_{2}\in\mathbb{R}^{3501}$ is the IR spectrum spanning 500–4000 cm-1 at 1 cm-1 resolution; and ${\bm{s}}_{3}\in\mathbb{R}^{3501}$ is the Raman spectrum on the same range. Together, these modalities provide a comprehensive description of the molecular signatures.
@@
Figure 1: (A) Overview of the DiffSpectra framework, illustrating the continuous-time forward diffusion process and reverse denoising process. The Diffusion Molecule Transformer (DMT) is used as the denoising network, and spectra encoded by SpecFormer serve as conditional input. (B) Architecture of the DMT, which processes node features, edge features, and atomic coordinates through three parallel streams with shared condition encoding, relational multi-head attention, and equivariant updates. (C) Architecture and pre-training strategy of SpecFormer, which encodes multi-modal spectra (UV-Vis, IR, and Raman) through a unified transformer encoder. It is pre-trained via masked patch reconstruction (MPR) and contrastive learning with 3D molecular structures.
@@
Within DiffSpectra, a continuous-time variational diffusion model progressively perturbs the molecular graph with Gaussian noise during the forward process, then denoises it in reverse, conditioned on spectra-derived information. A dedicated backbone, the Diffusion Molecule Transformer, parameterizes the denoising process while maintaining permutation and SE(3) equivariance. DMT is guided by spectral embeddings extracted by SpecFormer, a Transformer-based encoder pre-trained with both masked reconstruction and contrastive alignment to capture spectrum–structure correlations. This design allows the framework to achieve faithful and chemically meaningful structure elucidation, supported by physically grounded diffusion and rich spectral priors. More details of the full architecture and training objectives are presented in  Section 4.
@@
Table 1: Evaluation of basic metrics for molecular generation on the QM9S dataset. Metrics include stability, validity, uniqueness, and novelty, as well as distribution-based and 3D geometry-based metrics. The Train row reports statistics from the training set as a reference distribution. DiffSpectra is compared to unconditional molecular generative models.
Metric-2D Atom stable $\uparrow$ Mol stable $\uparrow$ V&C $\uparrow$ V&U $\uparrow$ V&U&N $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Train 99.9% 98.8% 98.9% 98.9% 0.0% 0.063 0.490 0.992 0.946 CDGS [36] 99.7% 95.1% 95.1% 93.6% 89.8% 0.798 0.493 0.973 0.784 JODO [37] 99.9% 98.8% 99.0% 96.0% 89.5% 0.138 0.522 0.986 0.934 DiffSpectra 99.9% 98.2% 98.4% 96.8% 93.7% 0.088 0.531 0.992 0.943
Metric-3D Atom stable $\uparrow$ Mol stable $\uparrow$ FCD $\downarrow$ Bond $\downarrow$ Angle $\downarrow$ Dihedral $\downarrow$ Train 99.4% 95.3% 0.877 $5.44\text{e}{-4}$ $4.65\text{e}{-4}$ $1.78\text{e}{-4}$ E-NF [38] 84.7% 4.5% 4.452 0.6165 0.4203 0.0056 G-SchNet [39] 95.7% 68.1% 2.386 0.3622 0.0727 0.0042 G-SphereNet [40] 67.2% 13.4% 6.659 0.1511 0.3537 0.0129 EDM [41] 98.6% 81.7% 1.285 0.1303 0.0182 6.64e-4 MDM [42] 99.2% 89.6% 4.861 0.2735 0.0660 0.0239 GeoLDM [43] 98.9% 89.7% 1.030 0.2404 0.0100 6.59e-4 JODO [37] 99.2% 93.4% 0.885 0.1475 0.0121 6.29e-4 DiffSpectra 99.2% 92.9% 0.943 0.1792 0.0054 2.01e-4

### v2

@@ Abstract
Molecular structure elucidation from spectra is a fundamental challenge in molecular science. Conventional approaches rely heavily on expert interpretation and lack scalability, while retrieval-based machine learning approaches remain constrained by limited reference libraries. Generative models offer a promising alternative, yet most adopt autoregressive architectures that overlook 3D geometry and struggle to integrate diverse spectral modalities. In this work, we present DiffSpectra, a generative framework that formulates molecular structure elucidation as a conditional generation process, directly inferring 2D and 3D molecular structures from multi-modal spectra using diffusion models. Its denoising network is parameterized by the Diffusion Molecule Transformer, an SE(3)-equivariant architecture for geometric modeling, conditioned by SpecFormer, a Transformer-based spectral encoder capturing multi-modal spectral dependencies. Extensive experiments demonstrate that DiffSpectra accurately elucidates molecular structures, achieving 40.76% top-1 and 99.49% top-10 accuracy. Its performance benefits substantially from 3D geometric modeling, SpecFormer pre-training, and multi-modal conditioning. To our knowledge, DiffSpectra is the first framework that unifies multi-modal spectral reasoning and joint 2D/3D generative modeling for de novo molecular structure elucidation.
@@
Molecular Structure Elucidation, Molecular Spectra, Diffusion Models, SE(3) Equivariance
@@ 1 Introduction
Molecular structure elucidation represents one of the fundamental challenges in chemistry, materials science, and biology, underpinning compound identification, reaction mechanism elucidation, and drug discovery [1, 2, 3]. This process involves the accurate determination of two-dimensional (2D) atomic connectivity and three-dimensional (3D) molecular conformation based on experimental data. Traditional approaches often rely on expert interpretation and substantial manual effort to propose and validate structural hypotheses. Although these methods have proven invaluable over decades of chemical research, the growing complexity of natural products, synthetic compounds, and pharmaceutical intermediates calls for more sophisticated computational approaches that can efficiently navigate the vast chemical space of possible structures.
@@
The integration of machine learning into spectroscopic analysis has opened new avenues for automating and accelerating molecular structure elucidation. Early applications of machine learning in molecular structure elucidation primarily adopted retrieval-based paradigms [1, 4, 5]. These approaches treat the task as a matching problem: given an experimental spectrum, the model searches a predefined molecular library to retrieve candidate structures whose spectra best align with that spectrum. Although efficient in practice, these methods are inherently constrained by the coverage and quality of the reference library. To mitigate this limitation, some efforts have proposed spectral prediction models, which simulate spectra from known molecular structures [6, 7, 8, 9, 10, 11, 12]. These predicted spectra are then used to expand the searchable library and improve retrieval coverage. However, such methods still rely on the existence of a finite reference library and thus fundamentally struggle to generalize to novel molecules outside it, thereby limiting their applicability in open-ended or exploratory chemical analyzes.
@@
To overcome reliance on predefined molecular libraries and advance toward genuinely data-driven elucidation, recent efforts have turned to more expressive neural architectures capable of inferring molecular structures de novo from spectral measurements. Initial attempts in this direction have framed the task as a predictive problem, training models to infer coarse structural features from spectra, such as molecular properties [13], fingerprints [14], fragments [15, 16], chemical formulas [17], or other approximate structural profiles [18, 19, 20, 21]. Although valuable, these coarse predictions fall short of achieving a complete molecular structure elucidation.
@@
In a shift toward complete structure elucidation, subsequent works have introduced generative models that take spectral data as input and produce discrete molecular representations, such as SMILES strings or molecular graphs. These models typically follow autoregressive architectures, generating molecular structures token-by-token using RNNs [22, 23, 24, 25] or Transformers [26, 16, 4, 27, 28, 29]. Despite their promise, these models face two fundamental limitations. First, generative models built on discrete SMILES strings or graphs inherently lack geometric inductive bias and cannot effectively capture 3D geometric information. This is critical because many spectral modalities, such as infrared (IR), Raman, and ultraviolet-visible (UV-Vis), are inherently dependent on the geometry of molecules and the associated energy states [30]. Second, all existing methods are limited to processing a single type of spectrum in isolation, whereas molecular structure elucidation, in practice, relies on the joint interpretation of multiple spectral modalities, each offering complementary structural cues [31]. Current models lack support for such multi-modal spectral reasoning, thereby constraining their accuracy and generalizability. These limitations highlight the need for a principled, unified generative framework that can integrate both geometric reasoning and multi-modal spectral conditioning.
@@
The inherent limitations of existing approaches motivate the development of more sophisticated generative models that can effectively bridge the gap between spectral observations and molecular structure determination. Diffusion models have recently attracted significant attention in generative modeling due to their ability to produce high-quality samples through iterative denoising processes [32, 33, 34, 35, 36]. This framework has demonstrated remarkable success in modeling complex, multi-modal distributions, generating diverse outputs, and showing particular promise in molecular applications [37, 3, 38, 39, 40, 41, 42, 43]. The ability of diffusion models to generate chemically meaningful structures and handle multi-modal distributions makes them particularly well-suited for molecular structure elucidation.
@@
In this work, we introduce DiffSpectra, a framework that leverages diffusion models to directly generate molecular structures from multi-modal spectra. Our approach addresses the structure elucidation problem as a conditional generation task, where spectral measurements serve as conditioning information that guides the iterative denoising process toward generating plausible molecular structures.
@@
The core of our framework is a joint 2D and 3D diffusion model that simultaneously generates molecular topological structures and geometric coordinates. Each molecule is represented as a geometric graph with node features (e.g., atom types), edge features (e.g., bond types), and 3D coordinates, thereby capturing both topological structure and geometric conformation in a unified representation. DiffSpectra formulates structure elucidation as a continuous-time conditional denoising process, where noise is progressively removed under spectral constraints. By jointly modeling topological and geometric components, the model ensures that the generated structures satisfy chemical bonding rules while preserving geometric consistency. This dual-space generation is essential for accurately elucidating molecules whose spectra reflect both bonding connectivities and 3D conformations, effectively leveraging the complementary strengths of topology and geometry to produce structurally consistent outputs.
@@
To parameterize the denoising process, we introduce the Diffusion Molecule Transformer (DMT), an SE(3)-equivariant architecture designed for molecular structure generation. DMT comprises a stack of transformer-style blocks operating on three parallel streams of node features, edge features, and coordinates, enabling rich interactions between topological and geometric representations. Through relational multi-head attention and equivariant geometric updates, DMT effectively captures spatial relationships while preserving chemical symmetry.
@@
To effectively condition the generation process on multi-modal spectra, we introduce SpecFormer, a Transformer-based spectrum encoder specifically designed to process and integrate multiple types of molecular spectra simultaneously, including IR, Raman, and UV-Vis spectra. SpecFormer partitions each spectrum into local patches, embeds them into a unified latent space, and employs multi-head self-attention to capture both intra- and inter-spectrum dependencies. To enhance the quality of spectral embeddings, we pre-train SpecFormer using masked patch reconstruction and contrastive alignment with molecular structures. The resulting spectral embeddings provide conditioning signals for DiffSpectra, effectively guiding the generation toward structures consistent with the measured spectra.
@@
We extensively evaluated DiffSpectra to assess its effectiveness in elucidating molecular structures from multi-modal spectra. Our results show that DiffSpectra generates chemically valid and stable molecular structures. When conditioned on spectra, our model achieves a top-1 exact-structure accuracy of 40.8%, and reconstructs key functional groups with over 96% similarity. We further show that using a pre-trained SpecFormer significantly improves structure-elucidation accuracy, and that combining multiple spectral modalities yields better performance than any single spectrum alone. Importantly, we find that sampling multiple candidates greatly enhances the hit rate: the top-10 accuracy rises to 99.49%, indicating that correct structures are consistently ranked among the plausible outputs. This highlights the practical value of diffusion-based sampling in structure elucidation, as generating a shortlist of candidates enables reliable downstream validation. Our ablations further confirm the advantage of generating 3D geometries and incorporating SE(3)-equivariance, and demonstrate that moderate sampling temperatures strike a good balance between diversity and accuracy. Further, DiffSpectra demonstrates robust generalization across molecules with varying numbers of atoms and exhibits an elucidation trajectory in which simpler substructures are resolved first, followed by more complex ones, as revealed by gradient analysis. These findings highlight the potential of spectra-guided diffusion models for accurate molecular structure elucidation.
@@ 2.1 Overview of DiffSpectra Framework
To achieve accurate elucidation of molecular structure from molecular spectra, we propose DiffSpectra, a spectra-conditioned diffusion framework that generates molecular structures guided by spectral information. The overall framework is illustrated in Figure 1.
@@
Formally, a molecule characterized by its structure and spectra is represented as $\mathcal{M}=(\mathcal{G},\mathcal{S})$. The molecular structure is modeled as a graph $\mathcal{G}=(\mathbf{H},\mathbf{A},\mathbf{X})$, where $\mathbf{H}\in\mathbb{R}^{N\times d_{1}}$ denotes node-level attributes (e.g., atom types and charges), $\mathbf{A}\in\mathbb{R}^{N\times N\times d_{2}}$ encodes pairwise edge features (e.g., bond existence and bond types), and $\mathbf{X}\in\mathbb{R}^{N\times 3}$ specifies the 3D coordinates of the atoms. Here, $N$ is the number of atoms, $d_{1}$ and $d_{2}$ are the node and edge feature dimensions. Atomic ordering is kept consistent across $\mathbf{H}$ and $\mathbf{X}$ to preserve the correspondence between atom types and spatial positions.
@@
Meanwhile, the molecular spectra are denoted as $\mathcal{S}=({\bm{s}}_{1},\ldots,{\bm{s}}_{|\mathcal{S}|})$, where $|\mathcal{S}|=3$ in our setting, corresponding to UV-Vis, IR, and Raman spectra. Specifically, ${\bm{s}}_{1}\in\mathbb{R}^{601}$ represents the UV-Vis spectrum from 1.5 to 13.5 eV with 0.02 eV intervals; ${\bm{s}}_{2}\in\mathbb{R}^{3501}$ denotes the IR spectrum spanning 500–4000 cm-1 at 1 cm-1 resolution; and ${\bm{s}}_{3}\in\mathbb{R}^{3501}$ denotes the Raman spectrum on the same range. Together, these modalities provide a comprehensive description of molecular signatures.
@@
Figure 1: Overview and architecture of the DiffSpectra framework. a, Schematic of the diffusion-based generative framework underlying DiffSpectra, comprising a continuous-time forward diffusion process and a reverse-time denoising process. The denoising network is implemented as the Diffusion Molecule Transformer (DMT), while spectral features encoded by SpecFormer provide conditional guidance along the denoising trajectory. b, Architecture of DMT, which jointly processes and denoises node features, edge features, and atomic coordinates through three parallel streams within a stack of equivariant blocks. The streams are interconnected by a relational multi-head attention mechanism and equivariant geometric updates, and are conditioned on shared spectral and time information. c, Architecture and pre-training strategy of SpecFormer, a unified Transformer encoder for multi-modal spectroscopic data (UV–Vis, IR, and Raman). SpecFormer is pre-trained with masked-patch reconstruction (MPR) and contrastive learning against corresponding 3D molecular structures.
@@
Within DiffSpectra, a continuous-time variational diffusion model progressively perturbs the molecular graph with Gaussian noise during the forward diffusion process, and then denoises it in reverse, conditioned on spectra-derived information. The denoising process is parameterized by the Diffusion Molecule Transformer, which preserves permutation and SE(3) equivariance. DMT is guided by spectral embeddings extracted by SpecFormer, a Transformer-based encoder pre-trained with both masked reconstruction and contrastive alignment to capture spectrum–structure correlations. This design enables faithful and chemically meaningful structure elucidation, supported by physically grounded diffusion and rich spectral priors. More details of the full architecture and training objectives are provided in Section 4.
@@
Table 1: Evaluation of basic molecular generation metrics on the QM9S dataset. Metrics include stability, validity, uniqueness, novelty, and distribution-based metrics. The Training Set row reports statistics of the training set as a reference distribution. DiffSpectra is compared against unconditional molecular generation models.
Models AtomStable $\uparrow$ MolStable $\uparrow$ V&U $\uparrow$ V&U&N $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Training Set 99.9% 98.8% 98.9% 0.0% 0.063 0.490 0.992 0.946 Unconditional Molecular Generation CDGS [44] 99.7% 95.1% 93.6% 89.8% 0.798 0.493 0.973 0.784 JODO [45] 99.9% 98.8% 96.0% 89.5% 0.138 0.522 0.986 0.934 Molecular Structure Elucidation DiffSpectra 99.9% 98.2% 96.8% 93.7% 0.088 0.531 0.992 0.943